OpenStudy (anonymous):

Here s my question: one angle is three times it's supplement increased by 20 deg. Find the measures of the two supplementary angles.

OpenStudy (rulnick):

x = 3 (180 - x + 20) = 3 (200 - x) = 600 - 3x So 4x = 600 x = 150 One angle is 150, the other is 30.

OpenStudy (rulnick):

Degrees, of course.

OpenStudy (anonymous):

x = 3 ( 180 -x ) + 20 x = 540 - 3x + 20 4x = 560 ->x = 560/4 = 140 degree its complement 40 degree